# Externally Valid Policy Evaluation Combining Trial and Observational Data

### v2

@@ Appendix A Proof
(*(*)( * ) $=\Bigg{[}\sum_{i\in\mathcal{S}_{+}}W_{i}\mathds{1}(L_{i}\leq\ell_{\alpha}^{\Gamma})\Bigg{]}\Bigg{[}\sum_{i\in\mathcal{D}^{\prime\prime}}\underline{W}_{i}^{\Gamma}\mathds{1}(L_{i}>\ell_{\alpha}^{\Gamma})+\overline{W}_{n+1}^{\Gamma}\Bigg{]}-\Bigg{[}\sum_{i\in\mathcal{D}^{\prime\prime}}\underline{W}_{i}^{\Gamma}\mathds{1}(L_{i}\leq\ell_{\alpha}^{\Gamma})\Bigg{]}\Bigg{[}\sum_{i\in\mathcal{S}_{+}}W_{i}\mathds{1}(L_{i}>\ell_{\alpha}^{\Gamma})\Bigg{]}$ (39)
$\geq\Bigg{[}\sum_{i\in\mathcal{D}^{\prime\prime}}W_{i}\mathds{1}(L_{i}\leq\ell_{\alpha}^{\Gamma})\Bigg{]}\Bigg{[}\sum_{i\in\mathcal{D}^{\prime\prime}}W_{i}\mathds{1}(L_{i}>\ell_{\alpha}^{\Gamma})+W_{n+1}\Bigg{]}-\Bigg{[}\sum_{i\in\mathcal{D}^{\prime\prime}}W_{i}\mathds{1}(L_{i}\leq\ell_{\alpha}^{\Gamma})\Bigg{]}\Bigg{[}\sum_{i\in\mathcal{D}^{\prime\prime}}W_{i}\mathds{1}(L_{i}>\ell_{\alpha}^{\Gamma})+W_{n+1}\Bigg{]}$ (40)
$=0.$ (41)
@@
We choose $\beta$ as in equation 16 to get the tightest limit. Note that in this context, $\mathbb{P}$ represents the probability over samples drawn from both $p(X,A,L|S=1)$ and $p_{\pi}(X,A,L|S=0)$. As $L_{n+1}$ is drawn from $p_{\pi}(X,A,L|S=0)$, we can express equation 43 as shown in equation 5 for the sake of notational clarity.
@@ Appendix B Numerical Experiments
Additional information of the numerical experiments outlined in Section 5 can be found in this supplementary material. All experiments were conducted using Version 1.7 of the Python implementation of xgboost. A comprehensive list of hyperparameters is available in Table 2 for the synthetic case and Table 3 for the NHANES case. The hyperparameters were selected through a random search involving 200 runs, employing 5-fold cross-validation with the F1 score as the optimization metric.
B.1 Reliability diagram
We use the reliability diagram technique to guide the lower bound of the range of miscalibration. We therefore need an estimate of the unknown odds
$\frac{p(S=0|X)}{p(S=1|X)},$ (44)
to be compare against the nominal odds
$\frac{\widehat{p}(S=0|X)}{\widehat{p}(S=1|X)}.$ (45)
Using equation 4, we have that
Take the expectation with respect to $X$, conditional on the nominal odds is in a specified interval (or bin) $I$, so that:
The expected odds is then estimated for each bin $I$ by counting samples from the target and trial distributions. For all numerical experiments, we use 5 bins.
@@ B.1 Reliability diagram
B.2 Synthetic data
We extend the experiments in Section 5.1 to include two extra target populations D, respective E. The full list of parameters used in equation 21 are given in Table 4.
Table 4: Means and variances of covariate distribution $p(X|S)$ in equation 21.
Population $\mu_{0,S}$ $\mu_{1,S}$ $\sigma_{0,S}^{2}$ $\sigma_{1,S}^{2}$
A $(S=0)$ 0.5 0.5 1.0 1.0
B $(S=0)$ 0.0 0.0 1.5 1.5
C $(S=0)$ 0.5 0.5 1.5 1.5
D $(S=0)$ 0.25 0.25 0.5 1.0
E $(S=0)$ 0.0 0.0 0.5 0.5
Trial $(S=1)$ 0.0 0.0 1.0 1.0
Figure 10 compares nominal sampling odds, $\text{odds}(X)$, obtained from the fitted models with the unknown odds, $\widehat{\text{odds}}(X)$, for the target populations C, D, respective E. In all these cases, the logistic model is misspecified and miscalibrated while the xgboost model provides odds that resemble the true ones.
@@ B.2 Synthetic data
Figure 10: True sampling odds, $\text{odds}(X)$, compared with nominal odds, $\widehat{\text{odds}}(X)$, obtained from logistic and xgboost models $\widehat{p}(S|X)$. The dots are a random subsample of the trial samples.
In Figure 11, we use the limit curves to evaluate the out-of-sample loss of a ‘treat all’ policy $\pi_{1}$, i.e.
$p_{\pi_{1}}(A=0\;|\;X)=1,$
for the target populations C, D, respective E. We start with population C in Figure 10(a). The limit curve derived for the baseline closely align with the logistic model when $\Gamma=1$, but is consistently slightly lower. For the certified curves the informativeness stays above 90% for odds miscalibration degrees $\Gamma\in[1,1.5]$. For the miscoverage gap, the baseline and the misspecified logistic model for $\Gamma=1$ are invalid, whereas the xgboost model has only a slight negative miscoverage gap. As the degree of miscalibration $\Gamma$ increases to 1.5, all limit curves indicate positive miscoverage gaps. In Figure 10(b), we evaluate the same policy $\pi_{1}$, but for population D. Now, the limit curves modelled using the baseline and logistic regression infer consistently higher losses than the curves modelled using xgboost when $\Gamma=1$. For the curve using the logistic model the informativeness stays above 90% for odds miscalibration degrees $\Gamma\in[1,1.5]$. The same figure for the xgboost model is 95%. All the miscoverage gaps are positive, but the curves held with xgboost are tighter. Finally, in Figure 10(c), we assess this for population E. The obtained limit curves are similar to the curves for population E but the informativeness stay above 95% for both models. Again, all the miscoverage gaps are positive, but the curves held with xgboost are tighter.
(a) Limit curve and miscoverage gap for target population C.
(b) Limit curve and miscoverage gap for target population D.
(c) Limit curve and miscoverage gap for target population E.
Figure 11: Evaluating ‘treat all’ policy $\pi_{1}$ for different target populations with degrees of miscalibration $\Gamma\in[1,1.5]$.
@@
In Figure 12, we use reliability diagrams to assess and compare the performance of logistic and xgboost models of the nominal odds for the target populations C, D, and E, respectively. For target population C, it is evident that the xgboost model shows a closer alignment with the diagonal compared to the logistic model. This discrepancy motivates a higher value for the parameter $\Gamma$ for the logistic model. This is in line with Figure 9(a) where the xgboost model is visually closer to the true model. For target populations D and E, both models deviate from the diagonal, although the xgboost model maintains a closer proximity to it when compared to the logistic model.
(a) Reliability diagram for target population C.
(b) Reliability diagram for target population D.
(c) Reliability diagram for target population E.
@@
Figure 12: Reliability diagram of the observed odds against the average predicted nominal odds obtained from logistic and xgboost-trained models.
In Figure 13, we compare the ‘treat all’ policy with a ‘treat none’ policy, i.e.,
$p_{\pi_{0}}(A=0\;|\;X)=1,$
for the target populations B, C, D and E. If their expected losses are estimated using equation 7, $\pi_{1}$ is the recommended choice over $\pi_{0}$ in all target populations. However, the tail losses certified for the ‘treat all’ policy $\pi_{0}$ are, specially for trial population B and C (Figure 12(a) and 12(b)), lower than the tail losses certified for $\pi_{1}$.
(a) Limit curves for target population B.
(b) Limit curves for target population C.
(c) Limit curves for target population D.
(d) Limit curves for target population E.
@@
Figure 13: Limit curves for $\pi_{0}$ and $\pi_{1}$ for different target populations certified for $\Gamma\in[1,1.5]$.
@@
For completeness, the actual degree of miscalibration for all target populations are visualized in Figure 14.
@@
(e) Degree of miscalibration for target population E.
Figure 14: The actual degree of miscalibration for different target populations. The dots are a random subsample of the trial samples, where the green ones corresponds to a degree of miscalibration within $\Gamma=1.5$ and the black ones to a degree of miscalibration not bounded by $\Gamma=1.5$.

### v3

@@ Appendix A Proof
()(*)=( ∗ ) = $\Bigg{[}\sum_{i\in\mathcal{S}_{+}}W_{i}\mathds{1}(L_{i}\leq\ell_{\alpha}^{\Gamma})\Bigg{]}\Bigg{[}\sum_{i\in\mathcal{D}^{\prime\prime}}\overline{W}_{i}^{\Gamma}\mathds{1}(L_{i}>\ell_{\alpha}^{\Gamma})+\overline{W}_{n+1}^{\Gamma}\Bigg{]}$ (37)
$-\Bigg{[}\sum_{i\in\mathcal{D}^{\prime\prime}}\underline{W}_{i}^{\Gamma}\mathds{1}(L_{i}\leq\ell_{\alpha}^{\Gamma})\Bigg{]}\Bigg{[}\sum_{i\in\mathcal{S}_{+}}W_{i}\mathds{1}(L_{i}>\ell_{\alpha}^{\Gamma})\Bigg{]}$ (38)
$\geq$ $\Bigg{[}\sum_{i\in\mathcal{D}^{\prime\prime}}W_{i}\mathds{1}(L_{i}\leq\ell_{\alpha}^{\Gamma})\Bigg{]}\Bigg{[}\sum_{i\in\mathcal{D}^{\prime\prime}}W_{i}\mathds{1}(L_{i}>\ell_{\alpha}^{\Gamma})+W_{n+1}\Bigg{]}$ (39)
$-\Bigg{[}\sum_{i\in\mathcal{D}^{\prime\prime}}W_{i}\mathds{1}(L_{i}\leq\ell_{\alpha}^{\Gamma})\Bigg{]}\Bigg{[}\sum_{i\in\mathcal{D}^{\prime\prime}}W_{i}\mathds{1}(L_{i}>\ell_{\alpha}^{\Gamma})+W_{n+1}\Bigg{]}$ (40)
[…]
@@
We choose $\beta$ as in equation 16 to get the tightest limit. As $L_{n+1}$ is drawn from $p_{\pi}(X,U,A,L|S=0)$, we can express equation 43 as shown in equation 4 for the sake of notational clarity.
@@ Appendix B Numerical Experiments
Additional information of the numerical experiments outlined in Section 5 can be found in this supplementary material and the code used for the experiments can be accessed here https://github.com/sofiaek/policy-evaluation-rct.
All experiments were conducted using Version 1.7 of the Python implementation of xgboost (Apache-2.0 License). A comprehensive list of hyperparameters is available in Table 2 for the synthetic case and Table 3 for the NHANES case. The hyperparameters were selected through a random search involving 200 runs, employing 5-fold cross-validation with the F1 score as the optimization metric. All experiments were performed on a laptop with the following specifications: Intel Core i7-8650 CPU @ 1.9GHz, 16 GB DDR4 RAM, and Windows 10 Pro 64-bit operating system. The experiments utilized only the CPU. The total time required to run all the experiments was approximately half an hour.
@@
B.1 Synthetic data
We extend the experiments in Section 5.1 to include two extra target populations C, respective D. The full list of parameters used in equation 20 are given in Table 4.
Table 4: Means and variances of covariate distribution $p(X,U|S)$ in equation 20.
Population $\mu_{0,S}$ $\mu_{1,S}$ $\mu_{U,S}$ $\sigma_{0,S}^{2}$ $\sigma_{1,S}^{2}$ $\sigma_{U,S}^{2}$
A $(S=0)$ 0.5 0.5 0.5 1.0 1.0 1.0
B $(S=0)$ 0.0 0.5 0.0 1.25 1.5 1.25
C $(S=0)$ 0.0 0.0 0.0 1.5 1.5 1.5
D $(S=0)$ 0.25 0.25 0.25 1.0 0.25 0.5
Trial $(S=1)$ 0.0 0.0 0.0 1.0 1.0 1.0
To illustrate the generality of the proposed methodology, we consider two different fitted models $\widehat{p}(S|X)$: a logistic model, which is conventionally used in the causal inference literature, and the more flexible tree-based ensemble model trained by xgboost. Figure 7 compares nominal sampling odds obtained from the fitted models with the unknown odds, $p(S=0|X)/p(S=1|X)$, for the target populations C, respective D. This corresponds to a case without unmeasured selection factors $U$. In all these cases, the logistic model is misspecified and miscalibrated while the xgboost model provides odds that resemble the true ones. Note that the proposed algorithm can handle any model if equation 3 is satisfied. However, a well-performing model is generally required to achieve a small $\Gamma$ and for a meaningful benchmark of the upper value of $\Gamma$.
@@ B.1 Synthetic data
Figure 7: Odds $p(S=0|X)/p(S=1|X)$ compared with nominal odds obtained from logistic and xgboost models $\widehat{p}(S|X)$. The dots are a random subsample of the trial samples.
@@
In Figure 8, we use the reliability diagram technique described in subsection 4.1 to assess and compare the performance of logistic and xgboost models of the nominal odds, $\widehat{\text{odds}}(X)$, for the synthetic data. For all numerical experiments, we use 5 bins. For target population A in Figure 8(a), both models are close to the diagonal and it is reasonable to believe that they are flexible enough to model the odds. For the other three populations, B, C and D, in Figure 8(b), 8(c) respective 8(d) it is evident that the xgboost model shows a closer alignment with the diagonal compared to the logistic model. This is in line with the results in Figure 5(b) and 7 where the xgboost model is visually closer to the true model.
(a) Reliability diagram for target population A.
(b) Reliability diagram for target population B.
(c) Reliability diagram for target population C.
(d) Reliability diagram for target population D.
@@
Figure 8: Reliability diagram of the observed odds against the average predicted nominal odds obtained from logistic and xgboost-trained models $\widehat{p}(S|X)$.
We now use the observed covariates to benchmark appropriate upper bounds for $\Gamma$ for all target populations as described in subsection 4.1. Figure 9(a) shows the evaluation with respect to population A. Assuming that the unmeasured selection factors $U$ are of similar strength as the weakest covariate, $X_{1}$, a $\Gamma$ value of $2.5$ could be a credible choice, as it covers $90\%$ of the odds ratios. Figure 9(b) shows the same evaluation with respect to population B. A $\Gamma$ value of $1.2$ could be a credible choice for the logistic regression model and a $\Gamma$ value of $2$ could be appropriate for the xgboost model. However, from Figure 5(b) we know that the logistic model is misspecified in this scenario. For population C in Figure 9(c), $1$ and $2.5$ seem to be suitable $\Gamma$ values for the logistic regression model respective the xgboost model. Similarly, for population D in Figure 9(d), $1.5$ and $2.2$ could be suitable values for the two models. The logistic model is again misspecified in populations C and D.
(a) Benchmarking for population A.
(b) Benchmarking for population B.
(c) Benchmarking for population C.
(d) Benchmarking for population D.
Figure 9: Benchmarking the degree of miscalibration $\Gamma$ using omitted covariates.
In Figure 10, we use the limit curves for the benchmark in equation 17 and the proposed method to evaluate the out-of-sample loss of the ‘treat all’ policy $\pi_{1}$, i.e.
$p_{\pi_{1}}(A=0\;|\;X)=1.$ (44)
Figure 10(a) shows the evaluation with respect to population A. When $\Gamma=1$ all the limit curves are similar. The curves for the proposed method also illustrate increasing the credibility of the models results in less informative inferences. However, their informativeness stays above 90% for odds miscalibration degrees $\Gamma\in[1,2]$. We also evaluate the miscoverage gap of the curves and observe that the benchmark and the limit curves for $\Gamma=1$ have a negative miscoverage gap. As the degree of miscalibration $\Gamma$ increases to $2$, the limit curves exhibit positive miscoverage gaps, where xgboost results in slightly less conservative inferences than the logistic model does. We continue with population B and C in Figure 10(b) respective 10(c). The limit curves derived for the baselines closely align with the curves modelled using logistic regression when $\Gamma=1$, but is consistently lower than the curves modelled using xgboost. For the certified curves the informativeness stays above 90% for odds miscalibration degrees $\Gamma\in[1,2]$. For the miscoverage gap, the baselines and the limit curves for $\Gamma=1$ are invalid. As the degree of miscalibration $\Gamma$ increases to $2$, all limit curves indicate positive miscoverage gaps. In Figure 10(d), we evaluate the same for population D. The limit curves modelled using the baseline and logistic regression when $\Gamma=1$ infer consistently higher losses than the curve modelled using xgboost. For the curve using the logistic model the informativeness stays above 90% for odds miscalibration degrees $\Gamma\in[1,2]$. The same figure for the xgboost model is approximately 95%. For the miscoverage gap, the baseline and the limit curves for $\Gamma=1$ are invalid. Again, as the degree of miscalibration $\Gamma$ increases to $2$, all limit curves indicate positive miscoverage gaps.
(a) Limit curve and miscoverage gap for target population A.
(b) Limit curve and miscoverage gap for target population B.
(c) Limit curve and miscoverage gap for target population C.
(d) Limit curve and miscoverage gap for target population D.
Figure 10: Evaluating ‘treat all’ policy $\pi_{1}$ for different target populations with degrees of miscalibration $\Gamma\in[1,2]$.
We now turn to comparing the ‘treat all’ policy with a ‘treat none’ policy, i.e.,
$p_{\pi_{0}}(A=0\;|\;X)=1,$ (45)
for population A. Their expected losses are estimated using equation 6 as $V^{\pi_{0}}_{\text{IPSW}}=1.64$ and $V^{\pi_{1}}_{\text{IPSW}}=1.54$, respectively. This evaluation suggests that $\pi_{1}$ is preferable to $\pi_{0}$. However, the limit curves presented in Figure 11(a) provide a more detailed picture in terms of out-of-sample losses: the tail losses certified for the ‘treat none’ policy $\pi_{0}$ are lower than those certified for $\pi_{1}$. This illustrates the cautionary principle built into the policy evaluations. Similar results are observed for target population D. While for target populations B and C, both $V^{\pi_{0}}_{\text{IPSW}}$ and $V^{\pi_{1}}_{\text{IPSW}}$ exhibit comparable sizes, the proposed limit curves still offer a more detailed understanding of out-of-sample losses.
(a) Limit curves for target population A.
(b) Limit curves for target population B.
(c) Limit curves for target population C.
(d) Limit curves for target population D.
@@
Figure 11: Limit curves for $\pi_{0}$ and $\pi_{1}$ for different target populations certified for $\Gamma\in[1,2]$.
@@
For completeness, the actual degree of miscalibration for all target populations are visualized in Figure 12 for the case without unmeasured selection factors $U$.
@@
Figure 12: The actual degree of miscalibration for different target populations. The dots are a random subsample of the trial samples, where the green ones correspond to a degree of miscalibration within $\Gamma=2$ and the black ones to a degree of miscalibration not bounded by $\Gamma=2$.
B.2 Seafood consumption policies
The National Health and Nutrition Examination Survey data (NHANES) is produced by federal agencies and is in the public domain, allowing it to be reproduced without permission. In our evaluation the data is split into observational data $\mathcal{D}_{0}$ and trial data $\mathcal{D}_{1}$ based on the covariates age, income, gender and smoking history (age and income are standardized), e.g.
$p(S=1|X,U)=0.25\cdot[f(X_{\text{age}})+f(X_{\text{income}})]+0.05\cdot[\mathds{1}(X_{\text{male}}=1)+\mathds{1}(X_{\text{smoking ever}}=1)]+0.3,$ (46)
where $f(\cdot)$ is the sigmoid function.